Police estimate that 77​% of drivers wear their seatbelts. They set up a safety​ roadblock, stopping cars to check for seatbelt use. If they stop 110 cars, what is the probability they find at least 18 drivers not wearing their​ seatbelts? Use a Normal approximation.

The probability is?

​(Type an integer or a decimal rounded to four decimal places as​ needed.)
Answer #1

We have here , n = 110 and p = 1-0.77= 0.23

n*p = 110* 0.23 = 25.3  

n*q = 110*(1-0.23) = 84.7

Therefore, we need to use here, Normal Approximation for the Binomial Distribution